Eric is making pizza dough. The recipe he is using will make 2 crusts and calls for ¼ cup of olive oil and 3/8 ounce of yeast.Eric gets distracted while making the dough, and accidentally puts 1 ounce of yeast into the bowl with the olive oil. How much more olive oil does Eric need to add so that the dough maintains the correct ratio of olive oil and yeast?
1
Expert's answer
2014-09-23T11:57:44-0400
Currently Eric has 1 / (3/8) = 8/3 times more yeast than usually. So, he needs to add (1/4)*(8/3) = 8/12 = 2/3 cup of olive oil to make excellent pizza dough as always.